OpenStudy (anonymous):

Please help fast I've been trying to complete this for an hour and I need to get it done before 6am The function f(t) = t2 + 4t − 14 represents a parabola. Part A: Rewrite the function in vertex form by completing the square. Show your work. (6 points) Part B: Determine the vertex and indicate whether it is a maximum or a minimum on the graph. How do you know? (2 points) Part C: Determine the axis of symmetry for f(t). (2 points)

OpenStudy (triciaal):

vertex form a(x-h)^2 + k h = -b/2a from standard form ax^2 + bx + c k = f(h) vertex (h, k)

OpenStudy (anonymous):

is that A?

OpenStudy (triciaal):

axis of symmetry is the line x = h

OpenStudy (anonymous):

Would this be a maximum because I don't see any negatives?

OpenStudy (triciaal):